# NeurStore: Efficient In-database Deep Learning Model Management System

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM International Conference on Management of Data; May 31–June 5, 2026; Bengaluru, Indiaisbn: 978-1-4503-XXXX-X/2026/06ccs: Information systems Database management system enginesccs: Information systems Database design and models
@@ 1. Introduction
Sectors such as finance (Hellerstein et al., 2012; Feng et al., 2012) and e-commerce (Park et al., 2022; Ooi et al., 2024) are rapidly adopting in-database AI-powered analytics in their critical business workflows, and with the advancements of AI, deep learning (DL) models have become prevalent. Consider purchase recommendations in e-commerce as an illustrative example, where items are recommended based on personal user profiles, including habits, occupations, and lifestyles. As user profiles typically contain sensitive information, such as salary and browsing history, in-database analytics is particularly suitable for handling such recommendation tasks. To enable precise and personalized recommendations, e-commerce vendors commonly deploy specialized DL models tailored to different users, regions, or customer segments. These specialized models are often derived from fundamental pre-trained DL models (Zeng et al., 2024), which may consist of dozens or even hundreds of layers, each potentially requiring gigabytes of storage (Su et al., 2024; Vartak, 2017; Miao et al., 2017). Further, as user profiles continuously evolve over time, these specialized models must be regularly updated or fine-tuned, leading to a steadily increasing number of models. As a result, efficient in-database DL model management (i.e., storing and loading DL models directly within DBMSs) has become a foundational capability for in-database AI-powered analytics.
@@
  • We implement NeurStore as a pluggable PostgreSQL extension and evaluate its performance against state-of-the-art systems. Experiment results demonstrate substantial gains in end-to-end AI-powered analytics performance, storage efficiency, and model saving and loading throughput.
  • We integrate NeurStore into DuckDB, and the performance evaluation confirms its general extensibility.
@@ 6.1.1. Workloads
To evaluate model management performance under realistic scenarios, we collect 800 DL models totaling 361GB, covering a diverse range of model architectures and sizes. Due to the space constraints, we provide a summary of these models in the extended version (Xiang et al., 2025). Before running the experiments, we perform a warmup phase to download all models from Huggingface (hug, 2025) and store them locally in advance. We feed the systems with read and write workloads, containing randomly selected models, to simulate model saving and loading operations. We first save the models into the evaluated system to measure write throughput and subsequently load them to measure read throughput.
@@ 6.2.1. End-to-end Performance Evaluation
We then evaluate the write and read throughput with varying numbers of clients ranging from 8 to 32. For write operations, the clients concurrently save models randomly chosen from the model pool described in Section 6.1.1. The results are shown in Figure 8(a). We can observe that the throughput of all systems increases as more clients are included due to the increased concurrency. NeurStore outperforms the baselines, achieving a peak throughput of 42 queries per minute compared with 37 for ELF and 17 for PostgresML. PostgresML performs the worst by storing the model with the TOAST mechanism in PostgreSQL. The models are divided into small chunks, which are fetched separately and reconstructed. This leads to high disk I/O overhead. ELF achieves throughput comparable to NeurStore across all numbers of clients, but remains behind due to higher I/O cost, especially when multiple models are stored concurrently. For read operations, each client continuously sends queries, with each query loading a random model. Figure 8(b) shows the read throughput. NeurStore outperforms ELF and PostgresML by up to $2.5\times$ and $1.4\times$, respectively. The improvement in the performance stems from three key designs. First, NeurStore maintains an in-memory index cache to avoid repeated fetching of base tensors, thus reducing the disk I/O. Second, NeurStore adopts an on-demand decompression strategy, where quantized deltas and base tensors are de-quantized at inference time, thereby eliminating the need for full model decompression prior to execution. Lastly, the flexible model loading enables NeurStore to load only the most significant 8 bits of the quantized deltas for inference, further reducing the disk I/O and memory bandwidth.
@@ 6.4.1. Performance Impact of Similarity Threshold
Figure 14. Performance of NeurStore on DuckDB
@@ 6.4.4. Performance Impact Across Models
We now evaluate the impact of flexible model loading and precision tolerance on model performance. In addition to the three tasks introduced in Section 6.1.1, we further include three more tasks to cover a wider range of model architectures: (1) summarization, where models generate abstractive summaries from dialogues, (2) object detection, where models identify multiple objects within images, and (3) tabular regression, where models predict continuous numeric values based on structured features. In total, our evaluation covers 200 models across six tasks. Due to the space constraints, we provide a summary of these models in the extended version (Xiang et al., 2025). We compress and decompress these models using a fixed precision tolerance of $5.96\times 10^{-8}$, and measure their absolute performance change. The results are shown in Figure 13. For ZFP, ELF, and NeurStore Full Loading, over 90% of models exhibit no performance change. For NeurStore Flexible Loading, 57 out of 200 models show a performance change less than 0.01%, and over 70% of models exhibit a change within 0.1%. More than 95% of models remain within a performance change of 1%. The increase in performance change is expected, since flexible loading restores only the most significant 8 bits of each delta tensor. Among all tasks, object detection models are most sensitive to precision loss, with an average performance change of 0.8%. This is because the object detection task requires predicting bounding boxes in images, where small changes in model weights can lead to high deviations in predicted object locations. On the contrary, image classification models show the smallest performance change, with an average of 0.009%, due to the simplicity of the task.
@@ 7. Related Works
There are several existing dedicated model management systems. ModelDB (Vartak, 2017) focuses on tracking model metadata, lineage, and experiment results to facilitate reproducibility and model governance, but it does not address model storage optimization. To reduce storage overhead, ModelHub (Miao et al., 2017) enables delta storage, which maintains the differences between fine-tuned and base models with explicit relations. However, ModelHub only captures pairwise differences between models and their predecessors, and does not account for the high entropy of floating-point weights, which limits its storage efficiency. In contrast, NeurStore targets tensor-level deduplication across the entire model collection, while incorporating a delta quantization algorithm that can efficiently compress high-entropy delta tensors.

### v2

@@ Abstract.
$c$s: Information systems Database management system enginesccs: Information systems Database design and models
@@ 1. Introduction
Sectors such as finance (Hellerstein et al., 2012; Feng et al., 2012) and e-commerce (Park et al., 2022; Ooi et al., 2024) are rapidly adopting in-database AI-powered analytics in their critical business workflows, and with the advancements of AI, deep learning (DL) models have become prevalent. Consider purchase recommendations in e-commerce as an illustrative example, where items are recommended based on personal user profiles, including habits, occupations, and lifestyles. As user profiles typically contain sensitive information, such as salary and browsing history, in-database analytics is particularly suitable for handling such recommendation tasks. To enable precise and personalized recommendations, e-commerce vendors commonly deploy specialized DL models tailored to different users, regions, or customer segments. These specialized models are often derived from fundamental pre-trained DL models (Zeng et al., 2024; Ooi et al., 2015; Luo et al., 2021; Gao et al., 2023), which may consist of dozens or even hundreds of layers, each potentially requiring gigabytes of storage (Su et al., 2024; Vartak, 2017; Miao et al., 2017). Further, as user profiles continuously evolve over time, these specialized models must be regularly updated or fine-tuned, leading to a steadily increasing number of models. As a result, efficient in-database DL model management (i.e., storing and loading DL models directly within DBMSs) has become a foundational capability for in-database AI-powered analytics.
@@
  • We implement NeurStore as a pluggable PostgreSQL extension and evaluate its performance against state-of-the-art systems. Experimental results demonstrate substantial gains in end-to-end AI-powered analytics performance, storage efficiency, and model saving and loading throughput. We further integrate NeurStore into two analytical databases (Raasveldt and Mühleisen, 2019; Schulze et al., 2024), and the performance evaluation confirms its general extensibility.
NeurStore is a component of NeurDB (Ooi et al., 2024; Zhao et al., 2025). It has been integrated into NeurDB as its native model storage.
The remainder of the paper is structured as follows. Section 2 provides the relevant background and presents the problem statement. Section 3 overviews the system architecture of NeurStore. Section 4 details the design of NeurStore. Section 5 describes the system implementation, and Section 6 presents the experimental results. Section 7 reviews the related works, and finally, Section 8 concludes the paper.
@@ 6.1.1. Workloads
To evaluate model management performance under realistic scenarios, we collect 800 DL models totaling 361GB, covering a diverse range of model architectures and sizes. Before running the experiments, we perform a warmup phase to download all models from Huggingface (hug, 2025) and store them locally in advance. We feed the systems with read and write workloads, containing randomly selected models, to simulate model saving and loading operations. We first save the models into the evaluated system to measure write throughput and subsequently load them to measure read throughput.
@@ 6.2.1. End-to-end Performance Evaluation
We then evaluate the write and read throughput with varying numbers of clients ranging from 8 to 32. For write operations, the clients concurrently save models randomly chosen from the model pool described in Section 6.1.1. The results are shown in Figure 8(a). We can observe that the throughput of all systems increases as more clients are included due to the increased concurrency. NeurStore outperforms the baselines, achieving a peak throughput of 39 queries per minute compared with 35 for ELF and 14 for PostgresML. PostgresML performs the worst by storing the model with the TOAST mechanism in PostgreSQL. The models are divided into small chunks, which are fetched separately and reconstructed. This leads to high disk I/O overhead. ELF achieves throughput comparable to NeurStore across all numbers of clients, but remains behind due to higher I/O cost, especially when multiple models are stored concurrently. For read operations, each client continuously sends queries, with each query loading a random model. Figure 8(b) shows the read throughput. NeurStore outperforms ELF and PostgresML by up to $2.5\times$ and $1.4\times$, respectively. The improvement in the performance stems from three key designs. First, NeurStore maintains an in-memory index cache to avoid repeated fetching of base tensors, thus reducing the disk I/O. Second, NeurStore adopts an on-demand decompression strategy, where quantized deltas and base tensors are de-quantized at inference time, thereby eliminating the need for full model decompression prior to execution. Lastly, the flexible model loading enables NeurStore to load only the most significant 8 bits of the quantized deltas for inference, further reducing the disk I/O and memory bandwidth.
@@ 6.4.1. Performance Impact of Similarity Threshold
Figure 14. Performance of NeurStore on DuckDB and ClickHouse
@@ 6.4.4. Performance Impact Across Models
We now evaluate the impact of flexible model loading and precision tolerance on model performance. In addition to the three tasks introduced in Section 6.1.1, we further include three more tasks to cover a wider range of model architectures: (1) summarization, where models generate abstractive summaries from dialogues, (2) object detection, where models identify multiple objects within images, and (3) tabular regression, where models predict continuous numeric values based on structured features. In total, our evaluation covers 200 models across six tasks. We compress and decompress these models using a fixed precision tolerance of $5.96\times 10^{-8}$, and measure their absolute performance change. The results are shown in Figure 13. For ZFP, ELF, and NeurStore Full Loading, over 90% of models exhibit no performance change. For NeurStore Flexible Loading, 57 out of 200 models show a performance change less than 0.01%, and over 70% of models exhibit a change within 0.1%. More than 95% of models remain within a performance change of 1%. The increase in performance change is expected, since flexible loading restores only the most significant 8 bits of each delta tensor. Among all tasks, object detection models are most sensitive to precision loss, with an average performance change of 0.8%. This is because the object detection task requires predicting bounding boxes in images, where small changes in model weights can lead to high deviations in predicted object locations. On the contrary, image classification models show the smallest performance change, with an average of 0.009%, due to the simplicity of the task.
@@ 7. Related Works
MLCask (Luo et al., 2021) is a model evolution management system that maintains the versioning of models in analytics pipelines, and it is complementary to NeurStore. Two earlier dedicated model management systems are ModelDB (Vartak, 2017) and ModelHub (Miao et al., 2017). ModelDB (Vartak, 2017) focuses on tracking model metadata and lineage to support machine learning workflow inspection, but it does not address model storage optimization. To reduce storage overhead, ModelHub (Miao et al., 2017) enables delta storage, which maintains the differences between fine-tuned and base models with explicit relations. However, ModelHub only captures pairwise differences between models and their predecessors, and does not account for the high entropy of floating-point weights, which limits its storage efficiency. In contrast, NeurStore targets tensor-level deduplication across the entire model collection, while incorporating a delta quantization algorithm that can efficiently compress high-entropy delta tensors.